OpenStudy (anonymous):

If Heather paid $692.75 in interest on a 6 year loan of $7,100. What was the interest rate?

OpenStudy (wolf1728):

Wow - that is one heck of a problem - but I think I can help. Principal - 7,100 Interest 692.75 Time 6 years Total paid on the loan = 7,100 + 692.75 = $7,792.75 7,792.75 / 72 = monthly payment = 108.2326388889 = $108.23 This calculator is a big help: http://www.1728.org/calcloan.htm Inputting 108.23 monthly payment 7,100 principal for 6 years equals 3.111 Interest rate Now we'll have to see how to calculate that (Unless you just need the answer then we're finished - but I think we'll need to do some more work).

OpenStudy (wolf1728):

Thanks for the medal - but there is more work to do.